### v2

# A Dynamic Model of Performative Human-ML Collaboration: Theory and Empirical Evidence

Abstract
Machine learning (ML) models are increasingly used in various applications, from recommendation systems in e-commerce to diagnosis prediction in healthcare. In this paper, we present a novel dynamic framework for thinking about the deployment of ML models in a performative, human-ML collaborative system. In our framework, the introduction of ML recommendations changes the data generating process of human decisions, which are only a proxy to the ground truth and which are then used to train future versions of the model. We show that this dynamic process in principle can converge to different stable points, i.e. where the ML model and the Human+ML system have the same performance. Some of these stable points are suboptimal with respect to the actual ground truth. We conduct an empirical user study with 1,408 participants to showcase this process. In the study, humans solve instances of the knapsack problem with the help of machine learning predictions. This is an ideal setting because we can see how ML models learn to imitate human decisions and how this learning process converges to a stable point. We find that for many levels of ML performance, humans can improve the ML predictions to dynamically reach an equilibrium performance that is around 92% of the maximum knapsack value. We also find that the equilibrium performance could be even higher if humans rationally followed the ML recommendations. Finally, we test whether monetary incentives can increase the quality of human decisions, but we fail to find any positive effect. Our results have practical implications for the deployment of ML models in contexts where human decisions may deviate from the indisputable ground truth.
1 Introduction
Human-ML collaboration is increasingly used in various applications, from content moderation in social media [1] to predicting diagnoses in healthcare [2, 3] and making hiring decisions in human resources [4]. Companies that implement human-ML collaborative systems face three crucial challenges: 1) ML models learn from past human decisions, which are often only an approximation to the ground truth (noisy labels); 2) ML models are rolled out to help future human decisions, affecting the data-generating process of human-ML collaboration that then influences future updates to ML models (performative predictions [5]); and 3) the quality of the human-ML collaborative prediction of the ground truth may change as a function of incentives and other human factors. These challenges create a dynamic learning process. Without access to the ground truth, it is often difficult to know whether the learning process will reach an equilibrium state with a good approximation of the ground truth, if it is interrupted at a sub-optimal level, or if it does not reach a stable state at all.
For intuition, we can focus on the decision of a healthcare company to develop and deploy an ML model to predict medical diagnoses from patient visits. The problem is made difficult by the fact that a doctor’s diagnoses can be wrong, and it is often too costly or time-consuming to identify the indisputable ground truth—i.e., the underlying true diagnosis of a patient—so the company typically uses all diagnoses to train their ML model, without distinction between good or bad diagnoses. In addition, the company typically evaluates the algorithm’s performance based on its ability to match those same doctor diagnoses, potentially replicating their mistakes. The dynamic deployment of updates to ML models that support doctor diagnoses could lead to a downward spiral of human+ML performance if the company deploys a bad model and the bad model adversely affects doctor decisions. Or, it can lead to continuous improvement until it reaches a stable point that is a good approximation to the indisputable ground truth. Without (potentially costly) efforts to measure the ground truth, the company has no way of distinguishing between downward spirals or continuous improvements.
This raises a multitude of empirical questions regarding the governing mechanisms of this dynamic system. How do humans improve on ML predictions of different quality levels, and do financial incentives matter? Will the dynamic learning process converge to a good equilibrium even without the company knowing the actual ground truth labels? How do we design an experiment to test such a system in an empirical context?
Contributions.
In this paper, we present a novel framework for thinking about ML deployment strategies in a performative, human-AI collaborative system. We present a simple theoretical framework to identify conditions under which ML deployment strategies converge to stable points that are a good approximation to the ground truth, and conditions under which there are downward spirals away from the ground truth. To validate our theory, we provide an empirical study in which humans solve knapsack problems with the help of machine learning predictions. We conducted a user study with 1,408 participants, each of whom solved 10 knapsack problems. Knapsack problems are particularly suited to explore our questions because we can train machine learning models to replicate human decisions, while having access to the underlying ground truth (the optimal solution) to evaluate the learning process. Additionally, knapsack problems can be hard for humans, making the task not obvious; and problem instances can be generated and labeled perfectly at negligible cost.
We highlight both theoretical and empirical contributions. Our theoretical framework introduces the collaborative characteristic function as the function mapping the performance of ML models with respect to the ground truth to the performance of humans using those models when making decisions (human+ML). We define utility the performance of decisions (made by ML alone, human alone, or human+ML) with respect to the ground truth. We note that utility cannot be easily quantified for many practical applications. Finally, we introduce the notion of collaborative learning paths, each of which characterizes a possible dynamic deployment strategy. We show conditions under which this dynamic system theoretically reaches a stable point of utility for the firm.
On the empirical side, we show how low-quality ML models can make humans perform worse than if they had no access to ML recommendations. Still, humans improve on ML models so that the deployment strategy of ML models with initial performance between 72% and 92% will converge to a performance that is around 92% of the value of the optimal knapsack solution. Two empirical findings constrain the system from converging to an even higher performance. First, humans do not respond to financial incentives for performance. Second, humans sometimes make decisions that are worse than the ML recommendation, despite the fact that it is fairly easy for them to compare their solution to the ML suggestion and pick the best of the two.
Our results have practical implications for the deployment of ML models when humans are influenced by those models but their decisions deviate from an unknown ground truth. First, performance metrics of ML models can be misleading when the learning objective is based on comparisons against human decisions and those decisions can be wrong. Companies should thus exert efforts to assess the quality of human decisions and take that into account when training ML models. For example, in the medical setting, human diagnoses should be first verified or confirmed by external experts, or patients should be followed up to confirm the validity of initial diagnoses. At a minimum, ML models should be trained on subsets of data for which there is enough confidence that the decisions are correct. Second, our work highlights the strategic importance of deploying ML models that allow for convergence to a stable point with higher utility than humans alone. Such convergence is not guaranteed and, as argued above, difficult to assess. Third, our work calls for the need to adopt a dynamic approach when deploying algorithms that interact with human decisions, and those interactions are used for future model building.
2 Related Work
There has been a growing body of work investigating various forms of human-AI collaboration. From learning-to-defer systems, where a model defers prediction tasks to humans if its own uncertainty is too high[6, 7, 8], to AI-assisted decision making where humans may or may not consult ML predictions to make a decision [9, 3, 2]. Several alternative decision mechanisms have also been explored [10, 11]. The application areas range from programming [12, 13], to healthcare [2, 3] and business consulting [14]. Related work also investigates factors influencing human-ML collaboration, such as explanations of ML predictions [15], monetary incentives [16], fairness constraints [17], and humans’ adaptability to model changes [18]. In this work, for the first time to the best of our knowledge, we examine the human+ML interaction from a dynamic perspective, where ML models learn from human decisions that are 1) the result of previous human+ML collaboration and 2) can arbitrarily deviate from the underlying ground truth.
This paper is also inspired by an extensive line of work on performative prediction [5, 19, 20, 21], a theoretical framework in which predictions influence the outcome they intend to predict. We adapt the ideas of performative prediction to a context of human-ML collaboration and extend it in three major ways: 1) In our setting, the model predictions change the quality of the human-ML labels as a proxy for the ground truth (e.g., a doctor diagnosis), but the ground truth is held constant (e.g., the true patient diagnosis); 2) We introduce the concept of utility, to quantify the quality of a solution with respect to the ground truth. There can be several stable points with respect to model parameters in the performative prediction framework, but not all of them are equally good at approximating the indisputable ground truth; and 3) The ground truth is unknown, and the mapping between human or ML labels and the ground truth is not fixed. To the best of our knowledge, we are the first to explore performative predictions where the company deploying ML models is unaware of the model’s performance relative to the ground truth, and only knows its similarity to human labels. Our empirical application is also novel in that it investigates the implications of performative predictions for human-ML collaboration.
3 Problem Statement
We consider a setting in which time is separable in discrete time epochs $t=1,...,T$. At each $t$, a firm deploys machine learning model $M_{t}\in\mathcal{M}$ with $M_{t}:\mathcal{X}\rightarrow\mathcal{Y}$. The model $M_{t}$ predicts a solution $Y\in\mathcal{Y}$ (e.g., a diagnosis) to a problem $X\in\mathcal{X}$ (e.g. the patient’s symptoms) as a function of past data. The firm employs expert humans $H\in\mathcal{H}$ with $H:\mathcal{X}\times\mathcal{Y}\rightarrow\mathcal{Y}$, who solve the problems with the help of the ML predictions. We will write $M_{t}(X)=Y_{M_{t}}$ and $H(X,Y_{M_{t}})=Y_{H_{t}}$. We assume that for all $X\in\mathcal{X}$, there exists an optimal solution $Y^{*}$, which is the undisputable ground truth.
The Firm’s Learning Objective.
In many real world applications, determining the ground truth label $Y^{*}$ can be extremely costly. For example, obtaining the correct medical diagnosis can often require the knowledge of various specialists (e.g., orthopedists, pediatricians, neurologists). Even when a single expert is enough, they can misdiagnose a patient’s symptoms. Yet, in many of these cases, using the human labels $Y_{H_{t}}$ as a proxy for $Y^{*}$ is the only feasible option to build ML models. We allow for the quality of $Y_{H_{t}}$ with respect to $Y^{*}$ to change. This means that two iterations of the ML model, $M_{t}$ and $M_{t+1}$, are trained on data from two different data generating processes, $(X,Y_{H_{t-1}})\sim~{}D_{t-1}$ and $(X,Y_{H_{t}})\sim D_{t}$, respectively.
Without access to $Y^{*}$, the only feasible learning objective for a firm that wants to update its model parameters at time $t$ is the comparison between the latest human-ML collaborative labels with the new predictions. [^1] We assume that models at time $t$ are trained exclusively on data from the previous period $t-1$, although we can generalize our setting to include any data points from 0 to $t-1$. For a given loss function $\mathit{l}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R_{+}}$ that is
$L(Y_{M_{t}},Y_{H_{t-1}}):=\underset{H\in\mathcal{H}}{\mathbb{E}}[\underset{(X,Y_{H_{t-1}})\sim D_{t-1}}{\mathbb{E}}\mathit{l}(Y_{M_{t}},Y_{H_{t-1}})].$ (1)
The firm wants to minimize the difference between the model predictions at time $t$ and the human labels at time $t-1$. We can write the firm’s problem as selecting a model $M_{t}$ to minimize the loss function in Equation 1:
$\underset{M_{t}\in\mathcal{M}}{minimize}\text{ }L(Y_{M_{t}},Y_{H_{t-1}}).$ (2)
For simplicity, we assume that at each time $t$, the firm collects enough data to perfectly learn the human-ML solution. In other words, with the optimal model, $L(Y_{M_{t}},Y_{H_{t-1}})=0$. We return to this assumption in Appendix A.7.
Utility.
In our scenario, the firm cannot quantify the true quality of a solution $Y$ with respect to $Y^{*}$. The loss in Equation 2 is just a surrogate for the loss $L(Y,Y^{*})$, which is impossible or too costly to obtain. The firm thus defines the human label as "ground truth," and maximizes the similarity between model and human solutions, without knowing how close the human or ML solutions are to the indisputable ground truth. In order to evaluate the firm’s progress in approximating $Y^{*}$, it is useful to define a measure of utility.
Definition 1.
(Utility) Let $d_{X}$ be a distance measure on $\mathcal{Y}$ with respect to a given $X\in\mathcal{X}$. The function $\mathbb{U}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a utility function on $\mathcal{X}\times\mathcal{Y}$, if $\forall X\in\mathcal{X},Y_{min},Y,Y^{\prime},Y^{*}\in\mathcal{Y}$
  1. 1.
    $\exists Y_{min}\in\mathcal{Y}:\mathbb{U}(X,Y)\in[\mathbb{U}(X,Y_{min}),\mathbb{U}(X,Y^{*})]$ (bounded)
  2. 2.
    $\exists\varepsilon>0:|d_{X}(Y,Y^{*})-d_{X}(Y^{\prime},Y^{*})|<\varepsilon\Rightarrow\mathbb{U}(X,Y)=\mathbb{U}(X,Y^{\prime})$ ($\varepsilon$-sensitive)
  3. 3.
    $d_{X}(Y,Y^{*})+\varepsilon<d_{X}(Y^{\prime},Y^{*})\Rightarrow\mathbb{U}(X,Y)>\mathbb{U}(X,Y^{\prime})$ (proximity measure)
The utility of a solution for the firm is maximal if $Y$ is $\varepsilon$-close to $Y^{*}$ with respect to the underlying problem $X$. The variable $\varepsilon$ should be interpreted as the threshold below which a firm perceives no difference between two outcomes, i.e., it does not care about infinitely small improvements.
Collaborative Characteristic Function.
As time $t$ increases, the firm hopes that the distributions $D_{t}$ shift closer to the optimal distribution $D^{*}$, where $(X,Y)=(X,Y^{*})$. In other words, for each model’s distance $d$, $d(D_{t},D^{*})>d(D_{t+1},D^{*})$. This could happen, for example, if humans were able to easily compare available solutions and pick the one that is closest to the indisputable ground truth. We can translate this continuous improvement into properties of the human decision function $H$ as follows: for all $t=1,...,T$ and $X\in\mathcal{X}$,
$\underset{H\in\mathcal{H}}{\mathbb{E}}[\mathbb{U}(X,H(X,Y_{M_{t}}))]=\mathbb{U}(X,Y_{M_{t}})+\delta_{M_{t}}.$ (3)
The firm’s hope is that $\delta_{M_{t}}\geq 0$ for $M_{t}$. Effectively, $\delta_{M_{t}}$ characterizes the human-ML collaboration for all utility levels of a model. If $\delta_{M_{t}}$ is positive, humans are able to improve on a ML prediction (and future model iterations will thus get better at approximating the ground truth). Instead, if $\delta_{M_{t}}$ is negative, humans will perform worse than the ML recommendations, and future model iterations will get progressively farther away from the ground truth. We define the function given by Equation 3 as the collaborative characteristic function:
Definition 2.
(Collaborative Characteristic Function) For a utility function $\mathbb{U}$, humans $H\in\mathcal{H}$, we define the collaborative characteristic function $\Delta_{\mathbb{U}}:\mathbb{R}\rightarrow\mathbb{R}$ as follows:
$\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))=\underset{H\in\mathcal{H}}{\mathbb{E}}[\mathbb{U}(X,H(X,Y_{M}))]=\mathbb{U}(X,Y_{M})+\delta_{M}.$
The function $\Delta_{\mathbb{U}}$ can take any arbitrary form. Several factors can affect $\Delta_{\mathbb{U}}$, e.g., monetary incentives and ML explanations. In subsequent sections, we empirically approximate one such function.
Collaborative Learning Path and Stable Points.
Although $\Delta_{\mathbb{U}}$ has infinite support, a firm will only experience a discrete set of utility values achieved by humans with the help of ML recommendations. We call this the collaborative learning path. It is characterized by $\Delta_{\mathbb{U}}$, the utility of the first deployed model $s$, and the number of deployment cycles $T$:
Definition 3.
(Collaborative Learning Path) Let $\Delta_{\mathbb{U}}$ be a collaborative characteristic function, $t=1,...,T\in\mathbb{N}_{\geq 1}$ the number of deployment cycles and $s=\mathbb{U}(X,Y_{M_{1}})$ the utility of the starting model. We define the collaborative learning path  to be the function
$\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)=\underset{H\in\mathcal{H}}{\mathbb{E}}[\underset{X\in\mathcal{X}}{\mathbb{E}}(\mathbb{U}(H(X,Y_{M_{t}}))].$
Definition 4.
(Stable Point) A stable point $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$ occurs at $t$ if for all $t^{\prime}\geq t$, $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t^{\prime})=\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$.
[…]
(a) Collaborative Improvement
(b) Collaborative Harm
Figure 1: Collaborative Improvement (left): The firm’s collaborative characteristic function  and one collaborative learning path, if humans improve on the ML solution. The x-axis denotes the model expected utility, the y-axis denotes expected human+ML utility. The firm deploys a first model with utility (s). Then humans use the model and improve utility by $\delta_{1}$, leading to expected human+ML utility (1). The firm learns a new model with utility (b) on the new data distribution. This is viable under the assumption that the new model has the same utility as the previous period’s human+ML labels, i.e., we can move horizontally from (1) to the 45-degree line at (b). Humans can further improve utility by $\delta_{2}$, which leads to expected utility (2). The dynamic improvement process continues until it reaches stable point utility (6-d). Collaborative Harm (right): The firm deploys a model with expected utility (s) but the humans, when interacting with the model, decrease utility by $\delta_{1}$, with expected utility (1). The firm will thus learn a model of utility (b) on the new distribution. The downward spiral continues until stable point (d).
We present the best-case and worst-case scenarios from Figure 1 as Propositions 1 and 2 below:
Proposition 1.
(Collaborative Improvement) If $\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))\geq\mathbb{U}(X,Y_{M})$ for all $M\in\mathcal{M},X\in\mathcal{X}$. Then $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$, is non-decreasing with $t=1,...,T$ and for sufficiently large $T$ it exists a $t^{\prime}\in[1,T]$ such that $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t^{\prime})$ is a stable point.
Proof.
(sketch) Because $\mathbb{U}$ is bounded, $\delta_{M}$ must be 0 in the extreme points. Furthermore, because of the $\varepsilon$-sensitivity of $\mathbb{U}$, the steps $t$ until reaching the maximum utility are also bounded. It follows that there exists a $t\in\mathbb{N}$ such that $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)-\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t+1)=0$, which is a stable point. See Appendix A.6 for the complete proof. ∎
Proposition 2.
(Collaborative Harm) If $\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))\leq\mathbb{U}(X,Y_{M})$ for all $M\in\mathcal{M},X\in\mathcal{X}$. Then $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$, is non-increasing with $t=1,...,T$ and for sufficiently large $T$ it exists a $t^{\prime}\in[1,T]$ such that $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t^{\prime})$ is a stable point.
Proof.
Similar to the proof of Proposition 1. ∎
In practice, a firm’s collaborative characteristic function can take any arbitrary shape, with portions above and portions below the 45-degree line. As long as the function is continuous, at least one stable point exists, and possibly more. When more than one stable point exist, the firm would like to reach the stable point with the highest utility (i.e., the highest point of the characteristic function lying on the 45-degree line). However, since the firm does not have access to the indisputable ground truth, when it reaches a stable point it does not know where such point lies on the 45-degree line.
In what follows, we empirically explore a context where 1) the ML model learns to imitate human decisions, and 2) it is easy for us to identify the indisputable ground truth. The setting allows us to draw at least a portion of the collaborative characteristic function, and explore the effects of human behavior on its shape, particularly the effect of monetary incentives and alternative selection criteria. We present study participants with instances of hard knapsack problems to answer the following research questions:
RQ1: How do monetary incentives affect human performance? To keep our treatment condition manageable, we explore the effect of different levels of performance bonuses on $\mathbb{U}(H(X,.))$, i.e., the human performance without ML recommendations.
RQ2: What is the shape of the human-ML collaborative characteristic function $\Delta_{\mathbb{U}}$? Here, we hold the performance bonus constant, and test humans’ effect $\delta_{M}$ on utility for different levels of ML performance. This will enable us to approximate $\Delta_{\mathbb{U}}$ for a specific task.
4 Experimental Setup
In this section, we describe our user study. The goal of our experimental setup is to simulate an environment in which users work on difficult tasks with the help of ML. The company responsible for deploying ML models does not know the optimal solution $Y^{*}$ (e.g., the true patient’s diagnosis), and it trains ML models to replicate experts’ decisions (doctor diagnoses). To evaluate how the company’s models perform against $Y^{*}$, we need a setting in which we, as researchers, know the quality of any solution Y using a utility function $\mathbb{U}(X,Y)$. This allows us to make absolute quality assessments of solutions. Note that this is often unattainable in practice, as we argued in the introduction. The Knapsack Problem is particularly well suited for this context.
The Knapsack Problem.
In our experiment, users solve instances of the knapsack problem. An instance involves selecting which of $n=18$ items to pack into a knapsack, each with a weight $w$ and a value $v$. The objective is to maximize value without exceeding the weight limit $W$ of the knapsack (between 5 and 250). We focus on the one-dimensional 0-1 knapsack problem, in which participants choose which items to pack (see Appendix 6 for a formal definition). We constrain the weights, values, and capacity of our instances to integer values, to make them easier to interpret by humans. We describe the details of the knapsack problem generation in Appendix A.10.
The knapsack problem has desirable properties for the empirical application of our dynamic framework. First, users do not require special training—beyond a short tutorial—to find a solution to the problem. Yet, the task is hard for humans, especially with a growing number of items [22]. Thus, the optimal solution $Y^{*}$ is not obvious. Second, we can generate solutions to the knapsack problem in two ways. The “optimal” solution can be found with dynamic programming. The “ML” solution can be found by imitating what humans select and computing the training loss as the difference between the items selected by participants versus items selected by a model.
This setup allows us to quantify the utility of the proposed solution relative to the optimal solution. We define utility for the knapsack problem as follows:
Definition 5.
(Economic Performance) For a knapsack instance $X=((w_{1},\cdots,w_{n}),(v_{1},\cdots,v_{n}),W)$ with optimal solution $\underset{x_{1},\cdot,x_{n};\sum_{i=1}^{n}x_{i}w_{i}\leq W}{\max}\sum_{i=1}^{n}x_{i}v_{i}=:Y^{*}$ and a valid solution $Y$ we call the function $\mathbb{U}_{\text{Econ}}(X,Y)=\frac{Y}{Y^{*}}$ the economic performance of $Y$ given $X$.
Appendix A.4 contains details about $\mathbb{U}_{\text{Econ}}(X,Y)$ and discusses our results using an alternative utility function. Note that there can be multiple optimal combinations of items to pack, but the optimal value $Y^{*}$ is unique.
Study Design.
We recruited participants from Prolific [^2] https://www.prolific.com/ exclusively from the UK to ensure familiarity with the currency and weight metrics used to describe the knapsack items and monetary incentives in the study. Appendix A.11 presents screenshots of the web interface for each step of the study. At the beginning of the study, participants received a tutorial on the knapsack problem, our web application’s interface, and the payment structure, described below. After the tutorial, the participants solved two practice problems and received feedback on their submission’s performance. For the main task, each participant received 10 knapsack problems generated by Algorithm 1. For each problem, they had 3 minutes to submit their solution. If the participant did not actively click on the submit button, the selected items were automatically submitted at the 3-minute mark. Participants could take unlimited breaks between problems. At the end of the study, we asked participants about their demographics, previous experience with the knapsack problem, and how much effort they put in solving the task.
Every participant received a base payment of £2.00 (approx. \$2.50) if they achieved at least 70% of the value of the optimal solution, averaged across the 10 knapsack instances they solved. We set the 70% threshold to discourage participants from randomly selecting items, as randomly-generated solutions that pick items until reaching the weight capacity have an average $\mathbb{U}_{\text{Econ}}$ around 60%.
Model None q1 q2 q3 q4 q5 q6
Mean $\mathbb{U}_{\text{Econ}}(X,Y)$ 0.717 0.800 0.844 0.884 0.899 0.920
SD . 0.083 0.105 0.098 0.105 0.088 0.085
No Bonus N=102
2-cent Bonus N=98
10-cent Bonus N=100+117 N=64 N=78 N=194 N=179 N=70 N=191
20-cent Bonus N=96
Table 1: Matrix of treatment conditions. The columns denote information on the ML recommendation performance. The rows denote bonus payments for performance. The number of study participants are presented in the relevant cells. We ran the 10-cent bonus treatment with no ML recommendation twice: once without a comprehension quiz for the bonus structure (100 participants) and once with the comprehension quiz (117).
Participants were randomly allocated into four monetary treatments and seven algorithmic recommendations (see Table 1). All monetary conditions were tested while users had no access to algorithmic recommendations. Participants in the No Bonus condition did not receive any additional payments beyond the base payment. Participants in the 2-cent Bonus condition received an additional £0.02 for each percentage point of $\mathbb{U}_{\text{Econ}}$ above 70%. For example, if a participant achieved on average $\mathbb{U}_{\text{Econ}}$ = 85%, they would receive £$\text{\textsterling}2.00+15\times\text{\textsterling}0.02=\text{\textsterling}2.30$. Participants in the 10-cent Bonus and 20-cent Bonus treatments had similar incentives for performance, but higher monetary rewards for each additional percentage point increase in performance (£0.10 and £0.20, respectively).
We ran the 10-cent Bonus treatment twice. In the second round, we introduced a comprehension quiz to ensure that our participants understood the payment structure. Within the 10-cent Bonus with bonus comprehension quiz, we randomized access to ML recommendations. Users were randomly allocated to one of seven ML treatments. The control group had no ML recommendations. The other six groups had access to recommendations from a progressively better ML model, as Table 1 shows on each of the last six columns. We picked ML models with varying degrees of performance to approximate the collaborative characteristic function from Figure 1. For details on the model training for the ML recommendations, see Appendix A.8.
5 Results
(a) Different bonus incentives.
(b) Different ML recommendations.
Figure 2: Human Performance Across Treatments. Error bars denote 95% confidence intervals based on standard errors clustered at the user level. Solid bars denote the average performance of the submitted solution, striped bars denote the performance if one picked the higher solution between the submitted solution and the provided ML recommendation.
A total of 1,408 participants completed the study; we removed 119 participants due to forbidden browser reloads or uses of the browser’s back-button, which left 1,289 for the analyses below. See Appendix A.9 for an overview of participants’ demographics. On average, participants’ compensation implied an hourly wage of £12.17 (\$15.22), which is above the UK minimum wage of £11.44. Please see Appendix A.3 for more payment details.
We start by discussing the null results of monetary performance incentives (RQ1). Figure 2(a) shows the results of our monetary incentive experiment. On average, user economic performance without any bonus was 89.7% (light blue bar). None of the bonus alternatives are statistically distinguishable from the control group, nor from each other, and their point estimates are all between 88.6% (for the 20-cent bonus) and 90% (for the 10-cent bonus).
The null effect of monetary incentives is not due to the fact that users did not understand the bonus structure. To test this hypothesis, we can compare the performance of users in the two 10-cent bonus treatments without algorithmic recommendations (third column in Figure 2(a) and first column in Figure 2(b), both yellow). These two treatments only differ by the fact that the one in Figure 2(b) had a comprehension quiz for the bonus structure. The difference in performance between the two treatments is a mere 0.9%, not statistically different from zero ($p=0.268$, based on standard errors clustered at the user level).
RQ2: Because monetary incentives do not change human performance, we test the introduction of ML recommendations with a single bonus structure, the 10-cent bonus. Figure 2(b) presents the results. Focusing on the solid bars, three insights are noteworthy. First, comparing the first two columns (yellow and blue), models with low economic performance seem to lead humans to perform slightly worse than if they were not supported by ML recommendations (89.4% versus 90.9%). This comparison is not statistically significant ($p=0.147$), likely due to low statistical power, but the level difference is not trivial. Despite this, humans do improve performance relative to the algorithmic recommendations (89.4% versus 71.8%, $p=1.8$e-$28$), a result we come back to in Section 5.1. This suggests that people might reduce their effort in solving the problem when they have access to recommendations, but they do not completely eliminate effort (Appendix Figure 18 shows no clear patterns in time spent per problem across treatment conditions). Second, models with better performance lead to increases in human performance, as evidenced by the progressively increasing performance from q1 (72%) to q6 (92%). Third, even if human performance increases with the performance of the ML recommendation, the increments in performance are quantitatively fairly small and sometimes statistically indistinguishable from one another, going from 89.4% when the model’s performance is 72%, to 92.6% when the model’s performance is 92%. [^3] Regression results, controlling for time taken to solve each problem, are presented in Appendix Table 3.
5.1 The Results Within Our Theoretical Framework
Figure 3: Empirical Collaborative Characteristic Function. Confidence intervals are based on standard errors clustered at the participant level.
Figure 3 embeds our empirical results in the theoretical framework presented in Section 3. On the x-axis, we plot the performance of the six ML models deployed in our study. On the y-axis, we plot the performance of the solutions submitted by humans who receive ML recommendations. Each of the points correspond to the six ML treatments of Figure 2(b). We linearly interpolate the estimated points to form an approximation of the collaborative characteristic function $\Delta_{\mathbb{U}}$ (solid blue line). The curve suggests that humans improve on the ML recommendations for ML performance levels between 70% and 92%. The estimated $\delta_{qi}$’s range from 17.5% ($p=1.8$e-$28$) for $q1$, to 0.5% ($p=0.46$) for $q6$. We denote $q6$ a stable point since the human improvement is estimated to be small and statistically indistinguishable from zero. The results imply that, for this portion of the domain, a firm can deploy a model with below-human performance and still converge to a stable point with 92% performance in subsequent deployments.
The $\delta_{qi}$ improvements are always positive (or indistinguishable from zero for q6), but they could have been even larger. Indeed, in this specific setting, as participants add items to the knapsack, in principle, they can easily compare the value of their solution to the value of the ML recommendation (both of which appear at the top of the interface, see Appendix Figure 15). If humans had picked the highest between their solution and the ML recommendation, the collaborative characteristic function would have shifted upward to the dashed green line in Figure 3, and the stable point would have achieved a higher performance than 92%. The discrepancy between the solid and dashed lines increases as the ML model improves, suggesting that even in a straightforward comparison, humans do not follow ML recommendations when it is in their best interest to do so (the difference can also be seen by comparing the solid and striped bars in Figure 2(b)). Appendix Figure 19 decomposes the net effect into two parts. On one hand, as the model performance improves, humans are more likely to follow its recommendations. On the other, when they do not follow the ML recommendation, as the model performance improves, it is much more likely that the submitted solution is inferior compared to the recommendation. Under both solid and dashed collaborative characteristic functions, we can imagine possible collaborative learning paths, $\mathbb{L}_{\Delta_{\mathbb{U}}}$. With this shape of $\Delta_{\mathbb{U}}$, the deployment decision is simple: all collaborative learning paths will eventually reach a stable point at above human performance.
6 Conclusions
We present a theoretical framework for human-ML collaboration in a dynamic setting where human labels can deviate from the indisputable ground truth. We introduce the collaborative characteristic function, which theoretically links the utility of ML models with respect to the indisputable ground truth, to the utility of humans using those same ML models to support their decisions. The collaborative characteristic function  allows for multiple collaborative learning paths, depending on the utility of the initially deployed ML model. Each of the collaborative learning paths characterizes a possible ML deployment strategy and its ensuing dynamic learning process. We theoretically show conditions under which this dynamic system reaches a stable point through dynamic utility improvement or deterioration. We then present the empirical results of a large user study, which allows us to estimate points on the collaborative characteristic function of the knapsack problem. For ML models of performance between 72% and 92%, our empirical results suggest that the collaborative characteristic function lies above the 45-degree line. Any collaborative learning path starting at utility between 72% and 92% will thus converge to a stable point with utility around 92%. We explore two factors that can shift the collaborative characteristic function. We find that monetary incentives do not seem to affect human performance. However, we find that wherever applicable, a simple post-processing step that picks the best among available solutions (as is possible for the knapsack problem) can substantially shift the collaborative characteristic function upward, leading to stable equilibria of higher utility.
Our work has a number of limitations. On the theoretical side, our collaborative learning paths assume that the firm is able to perfectly replicate human+ML performance in future ML models. Appendix A.7 discusses stability when learning does not exactly replicate previous human+ML performance. On the empirical side, to reduce costs while maintaining statistical power, we only randomized monetary incentives without ML recommendations, and we randomized the quality of ML recommendations while fixing monetary incentives. Studying the interaction of monetary incentives and ML performance is an important extension. The null result of monetary incentives should be interpreted within our context. First, the study participants received payments above minimum wage, and we only tested different levels of linear performance bonuses. It would be valuable to extend our work to evaluate the extent to which alternative base payments or non-linear bonuses may induce different levels of quality and effort by participants and thus collaborative characteristic functions of varying shapes. Our approximation of $\Delta_{\mathbb{U}}$ for the knapsack problem is naturally incomplete since we did not test every possible level of model performance. However, the six points of the curve that we empirically estimate make us fairly comfortable that a linear interpolation is reasonable, at least for model performances between our minimum and maximum.
Future work could investigate the properties of $\Delta_{\mathbb{U}}$ that guarantee a unique optimal stable point, both theoretically and empirically. Provided that researchers have access to the indisputable ground truth, further empirical investigations of collaborative characteristic functions could also shed light on the shape of those functions for practically relevant tasks such as medical diagnoses or hiring decisions. Future work should also discuss fairness aspects of this framework, e.g., whether or not fair stable points exist and how a firm can reach them. More generally, we hope this work generates more interest in studying settings where ML deployments lead to changes in the data generating process, which have broad managerial and practical applications.
Appendix A Appendix
A.1 Data and Code
The code for model training, data generation, the web application for user study and our data analysis and plotting can be found in retracted for anonymity; all files are part of the submission as .zip file
A.2 The Knapsack Problem
Definition 6.
[…]
A.3 Payment Details
We calculated the base payment assuming an average time of 19 minutes to complete the study. The base payment was adjusted upward if the median time to completion was longer than 19 minutes. We adjusted the payment despite the fact that many participants finished our survey but did not enter the completion code directly afterwards. This sometimes increased the median time to completion.
A.4 Analysis with Optimality
Definition 7.
(Optimality) For a knapsack instance $X$ with optimal solution $Y^{*}$ and a valid solution $Y$ we call the function $\mathbb{U}_{\text{Opt}}(X,Y)=\begin{cases}1&\text{if, }Y=Y^{*},\\ 0&\text{else }\end{cases}$ the optimality of $Y$ given $X$. Furthermore, we call $\underset{X}{\mathbb{E}}[\mathbb{U}_{\text{Opt}}(X,Y)]$, the optimal solution rate over all $X$.
Observation 1.
Economic performance and Optimality are utility functions (1).
Proof.
We start with the proof that Economic Performance is a utility function. 1) Economic performance is bounded between 0 (for an empty knapsack) and 1, for the optimal value of the knapsack. 2) There exists an $\varepsilon>0$, which is the minimum value of an item for the knapsack problem. The value of that item is the smallest possible distance between two solutions which are not equally good. 3) Because the $Y$ in our case is the sum of the values of the items in the knapsack and $Y*$ is the maximum possible value of the knapsack, any value that is closer to the optimal solution has also higher economic performance because the numerator grows. We chose $\varepsilon$ to be the minimum item value, thus this minimum increase in value between solutions is fulfilled. In summary, Economic Performance satisfies all three criteria of a utiltiy function.
We continue with the proof that optimality is a utility function. 1) it is 0 or 1 and thus bounded. 2) If we choose $10<\varepsilon<$ 1, then $\varepsilon$-sensitivity is satisfied. 3) Is always true for the choice of our $\varepsilon$. Assume for example $\varepsilon=0.5$, then it is that $d(1,1)+0.5<d(0,1)$ and $\mathbb{U}(1)>\mathbb{U}(0)$. This statement is true for all $10<\varepsilon<$ 1 which is what we specified for $\varepsilon$. ∎
Optimality is the function that indicates whether a solution to a knapsack problem has the optimal value or not. Figure 4 shows the empirical collaborative characteristic function for optimality as utility function. The humans achieve approximately $20\%$ optimalty without ML advice. The effect of human on human-ML performance is significant for all models ($p<0.001$). Interestingly, the effect is large even beyond human performance. Furthermore, for models q1,q2,3 with extremely low utility (average optimality of almost $0\%$), human effects on the overall outcome is large and close to human performance. As in Figure 3, the utility gain of rationally acting humans would have been larger for most models. Our observations suggest that stable points of optimality would lie above human performance without ML adivce.
Figure 4: Empirical Collaborative Characteristic Function for the "Optimality" utility function. Confidence intervals are based on standard errors clustered at the participant level.
A.5 Comments on the Definition of Utility
We want to denote that $\varepsilon$-sensitivity implies the following:
Observation 2.
$\exists\epsilon,\varepsilon>0:|d_{X}(Y,Y^{*})-d_{X}(Y^{\prime},Y^{*})|=\varepsilon\Rightarrow|\mathbb{U}(X,Y)-\mathbb{U}(X,Y^{\prime})|=\epsilon$
This means that there is a minimum utility change that we call $\epsilon$.
A.6 Proof of Propositon 1&2
Proposition 1
(Collaborative Improvement)
If $\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))\geq\mathbb{U}(X,Y_{M})$ for all $M\in\mathcal{M},X\in\mathcal{X}$. Then $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$, is non-decreasing with $t=1,...,T$ and for sufficiently large $T$ it exists a $t^{\prime}\in[1,T]$ such that $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t^{\prime})$ is a stable point.
Proof.
Let $t\in 1,...,T$ be the number of deployment (epochs) that a firm will make. The firm perfectly learns the data distribution in every epoch, in other words, we assume that $L(Y_{M_{t}},Y_{H_{t-1}}=0,\forall t$. Furthermore, it is $\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))\geq\mathbb{U}(X,Y_{M})$ for all $M\in\mathcal{M},X\in\mathcal{X}$.
We first show that $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$ is non-decreasing with $t$. For that, assume that there exists $t$ for which $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)>\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t+1)$. But $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t+1)=\underset{X\in\mathcal{X}}{\mathbb{E}}(\mathbb{U}(H(X,Y_{M_{t+1}})))\geq^{\delta_{i}\geq 0}\underset{X\in\mathcal{X}}{\mathbb{E}}(\mathbb{U}(Y_{M_{t+1}}))=^{L(Y_{M_{t+1}},Y_{H_{t}})=0}\underset{X\in\mathcal{X}}{\mathbb{E}}(\mathbb{U}(H(X,Y_{M_{t}})))=\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$. It follows that $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$ must be non-decreasing.
Now we show that there exists a $t^{\prime}\in[1,T]$ such that $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t^{\prime})$ is a stable point for sufficiently large $T$. For this, consider that $\mathbb{U}$ has a maximum $\mathbb{U}(Y^{*})$ (Property 1 (bounded) of definition 1) and there exists a minimum increment of utility $\epsilon$ (see A.5) in each deployment. If we do not achieve at least $\epsilon$ increment in utility, we have reached a stable point. Thus, we can write the maximum utility as $\mathbb{U}(Y^{*})=\mathbb{U}(Y_{M_{t}})+N\epsilon$. For sufficiently large ($T\geq N+1)$, this implies that we reached maximum utility with $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,T)$, and every deployment beyond that must have equal utility. ∎
Proposition 2
(Collaborative Harm)
If $\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))\leq\mathbb{U}(X,Y_{M})$ for all $M\in\mathcal{M},X\in\mathcal{X}$. Then $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t)$, is non-increasing with $t=1,...,T$ and for sufficiently large $T$ it exists a $t^{\prime}\in[1,T]$ such that $\mathbb{L}_{\Delta_{\mathbb{U}}}(s,t^{\prime})$ is a stable point.
Proof.
Analogous to the proof of Proposition 1. ∎
A.7 Perfect vs Imperfect learning
In this section we discuss what changes if we loosen the assumption $L(Y_{M_{t}},Y_{H_{t-1}})=0$. We call this assumption the "perfect learner" assumption because the firm perfectly learns the human labels from epoch $t-1$ with a model in epoch $t$. In the following, we consider an imperfect learner such that $L(Y_{M_{t}},Y_{H_{t-1}})=\sigma$. For now, we limit our discussion to the cases of Proposition 1 1 and Proposition 2 2.
In the case of collaborative improvement ($\Delta_{\mathbb{U}}(\mathbb{U}(X,Y_{M}))\geq\mathbb{U}(X,Y_{M})$) the human will improve on any model that the firm can deploy. However, if $L(Y_{M_{t}},Y_{H_{t-1}})=\sigma\Rightarrow(\mathbb{U}(Y_{M_{t}})-\mathbb{U}(Y_{M_{t-1}}))<0$ then our error $\sigma$ is larger than what we gain by putting a human in the loop. If the above statement is true for all $M_{t}$, then the imperfection has created collaborative harm and we arrived in the case of Proposition 2. However, this means that this would lead to a stable point. If we consider a second scenario where $L(Y_{M_{t}},Y_{H_{t-1}})=\sigma\Rightarrow(\mathbb{U}(Y_{M_{t}})-\mathbb{U}(Y_{M_{t-1}}))>0$ for all $M_{t}$. Then we are still in the scenario of collaborative improvement which means that we will reach a stable point. We can think about this when looking at Figure 1. An imperfect learner which is still a collaborative improvement setting is effectively like tilting the green dashed lines with a negative slope. We would still reach point 6-d but would require more iterations than with perfect learning.
In summary, this is an equally empirical question as it is for a perfect learner. The question is much humans improve the system’s performance and in the case of an imperfect learner how much of that improvement gets "eaten up" by learning errors.
A.8 Model Training
We release the code required for training our models, our model parameters and all predictions for the instances together with the instances that participants saw.
Learning to solve the knapsack problem is a research area for itself, however for the small, one-dimensional case of our experiment, it is possible on consumer hardware. We only train models for knapsack instances with 18 items. As input features we concatenate weights $w_{1},...,w_{18}$, values $v_{1},...,v_{18}$, the weight constraint $W$, the sum of the weights and the sum of the values. Thus, our input dimension is 39. Our goal was to train models with a broad spectrum of economic performances, not to solve the knapsack problem perfectly. We added 5 fully connected layers, 4 of them with ReLU activation functions. We use torch.Sigmoid() for our outputs. The output dimension was 18 and the output values in each index can be interpreted as the likelihood that the item belongs to a solution or not. For more details on the architecture, see our code.
We want to highlight two important aspects of how we thought about the model training. First, did not want to use any prior knowledge that a firm in our setting could not have either. For example, if we could have known the utility of a knapsack solution (economic performance or optimality) we could have just directly maximized it, or if we could know the optimal solution, we could have just used the distance to the optimal solution as our loss. Instead, we used the binary cross-entropy between the label and prediction as our loss. The label was a 18-dimensional 0-1 vector. If the i-th entry of this output vector is 1, it means that the i-th item is in the knapsack and otherwise not. Thus we simply minimized the differences between chosen items in our training data and those of our model. For us, this was a reasonable analogy for the application context of healthcare in which every "item" is a diagnosis or a symptom (e.g. an ICD10 code).
Because our financial budget was limited and we wanted to test multiple models, we trained all models on optimally solved knapsack instances. It would have also created a lot of overhead and space for errors if we would have collected the data of model q1 then trained q2 and rerun the user study. Training them all on generated labels made it possible to run more treatments at once. We still wanted to use ML models instead of solutions produced with dynamic programming, because we wanted to incorporate the distributional character of ML predictions (see Figure 8) and study the reaction to different quantiles of solution quality in greater detail in future work.
However, we had to include two pieces of prior knowledge in order to achieve better model performance (especially for q5 and q6). First, we sorted the items by density (value/weight). This is a big advantage in general, but only a small one for our knapsack instances because weights and values are strongly correlated. Second, we normalized weights and values in a pre-processing step. In our setting, both operations could not have been done by the firm (what is a normalized symptom)? However, with those minor modifications we were able to create a larger range of models without massive resources and still just immitate the "human" label without incorporating anything in the loss. In a post-processing step, we sorted the items by sigmoid outputs. We then added items to the knapsack until the weight constraint was reached. From that item selection, we calculated the actual knapsack values. For more details, please visit our github repository To be added after acceptance.
A.9 Overview statistics
Figure 5 shows the overview of the answer to the demographic questions in the end of our study. Most participants held an Undergraduate degree, were between 25 and 44 years old and have not heard about the knapsack problem before completing the study. $50.1\%$ of the participants identified as female $48.6\%$ as male and $0.8\%$ as non-binary or non-gender conforming. $96.8\%$ of the participants have not heard about the knapsack problem before this study. Figure 6 shows the perceived difficulty of the task for the participants, as well as the reported effort the participants put to complete the task. Most participants perceived the task as neutral to hard and put in large to very large effort (self-reportedly). Figure 7 shows how much effort people think they would have spent with or without the help of ML. It seems like participants who had no ML help think they would put less effort in the task. People who had the help of ML reported to put about as much effort as all participants reported to put in right now. Future work should investigate these perceptions in detail.
Figure 5: Highest level of education completed, age group, gender and whether participants have heard of the knapsack problem before this study.
Figure 6: Perceived difficulty of the task versus the reported level of effort participants reported in our study
Figure 7: How much effort participants thought that they would have spent with/without the help of ML
A.10 Generating Hard Knapsack Problems
[…]
Algorithm 1 Generate hard knapsack instance
number of items $n\geq 0$, knapsack capacity range $W_{min},W_{max}>0$
$W\leftarrow random.uniform.integer(W_{min},W_{max})$
$w\leftarrow random.uniform.integer(1,W,n)$triangleright n-dimensional vector of weights
$i\leftarrow 1$
while $i\leq n$ do
     $v_{i}\leftarrow max(1,random.uniform.integer(w_{i}-\lfloor\frac{W}{10}\rfloor,w_{i}+\lfloor\frac{W}{10}\rfloor)$
end while
A.11 Survey Design
Figure 8: Distribution of economic performances of solutions by the six models we deployed in our experiment.
Figure 9: Tutorial 1/5
Figure 10: Tutorial 2/5
Figure 11: Tutorial 3/5 (with ML treatment)
Figure 12: Tutorial 4/5 (with 10 cents/ppt monetary incentive)
Figure 13: Tutorial 5/5 (with comprehension quiz)
Figure 14: Feedback to a practice problem
Figure 15: Interface for the main task: 1) the knapsack capacity, 2) sum of weights of selected items, 3) sum of values of selected items, 4) remaining time, 5) items with weights and values, 6) machine learning solution (only visible if user receives corresponding treatment). Clicking on gray items adds them to the knapsack if the weight allows it, and clicking on green items removes them from the knapsack. The total weight and value of selected items is shown at the top and automatically updated.
Figure 16: Demographic questions after tasks
Figure 17: Score screen for performance feedback in the end
A.12 Futher Statistical insights
$\mathbb{U}_{\text{Econ}}(X,H(X))$ $\mathbb{U}_{\text{Econ}}(X,H(X))$
Intercept $0.7620^{***}$ $0.6957^{***}$
$(0.0184)$ $(0.0408)$
02-cent bonus $0.0003$ $0.1087^{*}$
$(0.0077)$ $(0.0484)$
10-Cent bonus $0.0011$ $0.0683$
$(0.0076)$ $(0.0499)$
20-Cent bonus $-0.0087$ $0.0787$
$(0.0079)$ $(0.0478)$
log(seconds spent) $0.0322^{***}$ $0.0481^{***}$
$(0.0039)$ $(0.0093)$
02-cent bonus $\cdot$ log(seconds spent) $-0.0260^{*}$
$(0.0112)$
10-cent bonus $\cdot$ log(seconds spent) $-0.0161$
$(0.0115)$
20-cent bonus $\cdot$ log(seconds spent) $-0.0210$
$(0.0113)$
N 3,960 3,960
Adj.R2 0.0613 0.0661
$\mathbb{U}_{\text{Econ}}(X,H(X,Y))$ $\mathbb{U}_{\text{Opt}}(X,H(X,Y))$
Intercept $0.8082^{***}$ $-0.0297$
$(0.0049)$ $(0.0245)$
q1 ($72\%$) $-0.0131$ $-0.0408$
$(0.0044)^{**}$ $(0.0260)$
q2 ($80\%$) $0.0011$ $-0.0161$
$(0.0043)$ $(0.0283)$
q3 ($84\%$) $0.0048$ $-0.0324$
$(0.0049)$ $(0.0198)$
q4 ($88\%$) $0.0151^{***}$ $0.0333$
$(0.0033)$ $(0.0214)$
q5 ($90\%$) $0.0247^{***}$ $0.0518^{*}$
$(0.0043)$ $(0.0263)$
q6 ($92\%$) $0.0211^{***}$ $0.0880^{***}$
$(0.0033)$ $(0.0213)$
log(seconds spent) $0.0240^{***}$ $0.0533^{***}$
$(0.0010)$ $(0.0045)$
N 8,930 8,930
Adj.R2 0.0733 0.0281
Table 2: Linear regressions with clustered standard errors on participant id. Effect of monetary incentive on $\mathbb{U}_{\text{Econ}}$ of human solutions (left). Effect of ML recommendation on different levels of economic performance $\mathbb{U}_{\text{Econ}}$ (right). Standard errors in parentheses. * $p<0.05$, ** $p<0.01$, *** $p<0.001$.
(a) Different bonus incentives.
(b) Different ML recommendations.
Figure 18: Time Spent Across Treatment Conditions
$\mathbb{U}_{\text{Econ}}(X,H(X))$
(1) (2) (3)
Intercept $0.7620^{***}$ $0.7676^{***}$ $0.8082^{***}$
$(0.0184)$ $(0.0276)$ $(0.0049)$
2-cent Bonus $0.0003$
$(0.0077)$
10-cent Bonus $0.0011$
$(0.0076)$
20-cent Bonus $-0.0087$
$(0.0079)$
Comprehension Quiz $0.0104$
$(0.0076)$
q1 ($72\%$) $-0.0131$
$(0.0044)^{**}$
q2 ($80\%$) $0.0011$
$(0.0043)$
q3 ($84\%$) $0.0048$
$(0.0049)$
q4 ($88\%$) $0.0151^{***}$
$(0.0033)$
q5 ($90\%$) $0.0247^{***}$
$(0.0043)$
q6 ($92\%$) $0.0211^{***}$
$(0.0033)$
log(seconds spent) $0.0322^{***}$ $0.0317^{***}$ $0.0240^{***}$
$(0.0039)$ $0.0064$ $(0.0010)$
N 3,960 2170 8,930
Adj.R2 0.0613 0.0506 0.00733
Included Bonus Treatments All 10-cent 10-cent
Included ML Treatments No ML No ML All ML
Comprehension Quiz No Both Yes
Table 3: Linear regressions of economic performance $\mathbb{U}_{\text{Econ}}$ on dummies for the various treatment conditions. Column 1 includes all treatment conditions without ML recommendations and without comprehension quiz. It tests the difference in performance across different bonus levels. Column 2 includes the two treatment conditions without ML recommendation and with 10-cent bonus. The difference between the two treatment conditions is the presence of a comprehension quiz for the bonus structure. Column 3 includes all treatments with comprehension quiz and 10-cent bonus. It tests the difference in performance across ML recommendations with different performance. Standard errors, in parentheses, are clustered at the participant level. * $p<0.05$, ** $p<0.01$, *** $p<0.001$.
(a) Rate of ML advice usage increased with better performance.
(b) Participants ignore the ML recommendation with better performance.
Figure 19: ML-usage increased with better ML performance. Share of ignored ML solutions did also increase with better performance.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All claims are supported by our theory in section 3 and/or by empirical results, presented in section 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We included the limitations in the conclusion of the paper (section 6). We address limitations regarding experimental design choices, statistical power, theoretical assumptions, and the limitations of our interpretations. In section 6, we also acknowledge that future work can substantially expand on this paper, by for example addressing relevant fairness aspects of our novel framework (beyond the scope of this work) and conducting additional empirical studies.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide a full set of assumptions and sketches of the proofs of Propositions 1 and 2 in section 3. A complete and correct proof for both propositions is provided in Appendix A.6.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We disclose and release the complete setup (section 4, including description of treatments, payments, and crowd-sourcing platform) and results (section 5) of our experiments. Appendices and supplementary files provide further details: the code for the web-app on which participants solved the knapsack problems and answered the survey, the data of all experiments, the code for model training, data generation, data analysis, and hardware requirements. Every step of our study is reproducible with a new set of participants. We also provide screenshots of the web app and a description of the experimental design and task.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide open access to data and code for everything we did with sufficient comments to reproduce our experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe essential information in the main paper (section 4). Model training details are given in the appendix and the released code.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The error bars are all defined in the relevant figure captions, as well as when discussing results in section 5. We also report p-values where appropriate.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report hardware requirements in the experimental details in the appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: All participants received appropriate compensation for their efforts. The design of our study conforms to the highest standards of privacy. Participants give their consent to participate in our study after receiving information about all the data that we collected. We discuss broader impacts of our work in the conclusion of the main paper.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss potential broader impacts of this work at the end of our introduction (section 1) and in our conclusions (section 6).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The work and models released in this paper do not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite work that is not ours wherever we use it. The data we collect are all primary and original data.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release our experimental code with comments and explanations that make it easily reproducible. We describe our experiments in detail in the main paper and appendix. Code and data for any of the analyses performed (human study, model training, data analysis) will be uploaded as a zip file and released on Github upon acceptance.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: We include the code for the experimental platform, the instructions and interface as screenshots. We report the compensation of the participants in the main text (sections 4 and 5) and in the appendix.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: An IRB approval was not required from the institutions of the authors for this study. The study has been registered and documented in a data protection registry of the main author’s institution and has been approved by the data privacy officer of their institution. The consent form and imprint of the experimental platform contained contact information from the main author. It also included information about third-parties in case participants wanted to contact them with concerns about data privacy and research ethics.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.